# FoMo: A Foundation Model for Mobile Traffic Forecasting with Diffusion Model

### v1

@@ Abstract.
Mobile traffic forecasting allows operators to anticipate network dynamics and performance in advance, offering substantial potential for enhancing service quality and improving user experience. However, existing models are often task-oriented and are trained with tailored data, which limits their effectiveness in diverse mobile network tasks of Base Station (BS) deployment, resource allocation, energy optimization, etc., and hinders generalization across different urban environments. Foundation models have made remarkable strides across various domains of NLP and CV due to their multi-tasking adaption and zero/few-shot learning capabilities. In this paper, we propose an innovative Foundation model for Mobile traffic forecasting (FoMo), aiming to handle diverse forecasting tasks of short/long-term predictions and distribution generation across multiple cities to support network planning and optimization. FoMo combines diffusion models and transformers, where various spatio-temporal masks are proposed to enable FoMo to learn intrinsic features of different tasks, and a contrastive learning strategy is developed to capture the correlations between mobile traffic and urban contexts, thereby improving its transfer learning capability. Extensive experiments on 9 real-world datasets demonstrate that FoMo outperforms current models concerning diverse forecasting tasks and zero/few-shot learning, showcasing a strong universality. We further deploy the FoMo on the JiuTian optimization platform of China Mobile, where we use the predicted mobile data to formulate network planning and optimization applications, including BS deployment, resource block scheduling, and BS sleep control.
Mobile traffic forecasting; foundation model; diffusion model; spatio-temporal prediction
@@ 1. Introduction
In recent years, foundation models (Brown, 2020; Touvron et al., 2023; Radford et al., 2021) have made substantial strides in natural language processing and computer vision. These models are reshaping the AI ecosystem by harnessing their powerful data processing, generalization, and zero/few-shot learning capabilities. An increasing number of specialized domains have developed foundational models tailored to their specific data and contextual demands, including healthcare, medicine, urban navigation, intelligent manufacturing, and beyond (Yuan et al., 2024b; Yeh et al., 2023; Kang, 2024; Bozic et al., 2024). Mobile networks and computing encompass massive amounts of mobile traffic, user, and geographical data, providing natural data support for foundation models. However, such dedicated foundation models for mobile network domains have yet to be established. We hence aim to construct a foundation model for mobile traffic forecasting, which can handle diverse features of large-scale mobile data while retaining the generalization needed across multiple cities and scenarios (Wu et al., 2021b; Chai et al., 2024a; Yin et al., 2022).
@@
Constructing such a foundation model of mobile traffic forecasting is vital for mobile networks (Feng et al., 2024; Wu et al., 2021a; Zhou et al., 2023). On the one hand, mobile traffic forecasting offers great potential for network planning and optimization. It enables operators to anticipate traffic dynamics, facilitating proactive perceptions of resource utilization and service quality, and allowing for the preemptive development of optimization strategies. On the other hand, mobile networks encompass a variety of optimization tasks, including radio resource scheduling (Miao et al., 2012; Lohani et al., 2016), Base Station (BS) deployment (Dong et al., 2020; Fahim and Gadallah, 2023), and antenna configuration (Mei and Zhang, 2023; Lin et al., 2022), etc. These tasks involve diverse objectives like throughput, coverage, and energy efficiency, which impose distinct requirements on traffic forecasting. For example, radio resource scheduling prioritizes short-term traffic dynamics to improve user experience (Chen et al., 2020b; Ghandri et al., 2024), whereas BS deployment focuses on long-term traffic patterns within a region to match network demands (Zhao et al., 2021; Wang et al., 2021). Regarding the differentiated forecasting needs of optimization tasks, current methods typically employ one-to-one approaches: designing customized models by leveraging task-specific data (Yang et al., 2019; Hu et al., 2023; Zhou et al., 2023; Li et al., 2024), and focus on one single application scenario, making it challenging to generalize the models across different cities and tasks.
On the contrary, foundation models possess large-scale learnable parameters with powerful feature representation and generalization capabilities, enabling them to learn universal features from vast amounts of mobile traffic data (Liang et al., 2024; Chaccour et al., 2024). Moreover, it is endowed with robust few/zero-shot abilities to be applied across different cities and scenarios (Cai et al., 2023; Rahman et al., 2018). By constructing foundation models for traffic forecasting, it becomes possible to achieve precise traffic predictions for various scenarios and cities via one unified model, thereby efficiently addressing the diverse tasks involved in mobile networks.
However, designing the foundation model remains a significant challenge. Although numerous notable works have emerged in the area of mobile traffic forecasting (Gong et al., 2023; Kang et al., 2024; Pandey et al., 2024; Kavehmadavani et al., 2024; Xu et al., 2021; Zhang et al., 2023), the majority of these models are typically tailored to specific tasks that fall short of the universality expected from foundation models. Specifically, current mobile traffic forecasting models face three key limitations:
i) Limited generalization. Mobile traffic data is inherently shaped by the spatio-temporal dynamics of population distribution and communication demands. Due to variations in geographic environments, lifestyle habits, and urban layouts across different cities, mobile traffic can differ significantly (Yin et al., 2020; Xu et al., 2017). With relatively small parameters, current models struggle to capture the diverse spatio-temporal patterns inherent in large-scale data across multiple cities. Additionally, it is challenging to encapsulate the complex correlations between contextual factors and mobile traffic, resulting in poor transferability in multi-city scenarios.
ii) Constrained adaptability to multiple scenarios. Mobile traffic forecasting is extensively applied across varying optimization scenarios. However, current models are often designed with specialized modules tailored to specific tasks. For instance, in short-term forecasting, models usually focus on capturing traffic fluctuations that employ autoregressive or event-driven methods. In contrast, long-term predictions emphasize the regular patterns of traffic and typically utilize time series decomposition techniques. These dedicated models increase design complexity and raise deployment costs when applied to diverse scenarios.
@@
iii) Lack of universal representation of mobile traffic. Unlike standardized formats in language and vision domains, mobile traffic data is inherently heterogeneous of various collection granularity and scope. For example, Measure Report (MR) data primarily collects millisecond-level user traffic, while Performance Management (PM) data gathers cell-level traffic statistics over 15-minute intervals (Naboulsi et al., 2016; Wang et al., 2017b), leading to the absence of a unified representation akin to that found in natural language. Consequently, it is challenging to directly apply pre-trained models from the natural language/visual domains to mobile traffic data. Although some efforts have been made to reprogram mobile traffic data into a natural language format (Gruver et al., 2023; Jin et al., 2024), this approach heavily relies on the quality of manually crafted prompts, making it difficult to capture a universal representation of mobile traffic.
@@
To tackle the challenges, we propose a Foundation model for Mobile traffic forecasting (FoMo), which aims to learn universal features of mobile traffic data and to handle multitasks in mobile networks across multiple cities, thereby establishing a one-for-all forecasting model. First, inspired by Sora (Brooks et al., 2024), FoMo adopts the transformer-based diffusion model as the backbone instead of the U-Net structure, to help the model understand the diverse features of massive mobile data. Furthermore, We propose a contrastive diffusion algorithm and modify the variational lower bound of the training objective by analyzing the cross-entropy between mobile traffic and contextual features, enabling the model to better integrate environmental information, thereby enhancing generalization capabilities and addressing the first research challenge. Second, we adopt a mask-based and self-supervised training paradigm (He et al., 2022; Wei et al., 2023), where we categorize traffic forecasting in mobile networks into three tasks: short-term prediction, long-term prediction, and generation. We design the corresponding masking strategies to enable the model to learn data features for various tasks and adapt to multiple tasks, thus addressing the second research challenge. Finally, we develop a mobile tokenization scheme to represent mobile traffic. The token standardizes mobile traffic data collected from various periods and diverse urban regions into unified units. The model leverages these tokens to learn the underlying patterns of large-scale mobile traffic data, enabling more accurate forecasting.
@@
$\bullet$ To the best of our knowledge, it is the first foundation model designed for mobile traffic forecasting. The proposed model enables various forecasting tasks in mobile networks across different urban environments under a unified framework, assisting network operators in achieving highly efficient network planning and optimization.
@@
$\bullet$ We develop our foundation model using a masked diffusion approach with spatio-temporal masking strategies tailored for diverse forecasting tasks, including short/long-term predictions and distribution generation. To strengthen the correlation between contextual features and mobile traffic, we further propose a context-aware contrastive learning fine-tuning strategy, which can enhance forecasting and transfer learning capabilities.
$\bullet$ We conduct extensive experiments with 4G and 5G mobile traffic data from 9 real-world datasets. The experimental results validate FoMo’s superior transferability, multi-task capabilities, and robust few/zero-shot performance in unseen scenarios. We further experiment with the scaling properties of our model, preliminarily uncovering the scaling law of the mobile traffic foundation model concerning data size, model scale, and model performance.
2. Preliminaries and Problem Formulation
2.1. Diffusion model and training object
Denoising Diffusion Probabilistic Model (DDPM). The underlying idea of DDPM is to use two Markov chains to characterize the transition from original data to noise data (Mukhopadhyay et al., 2023). The forward chain gradually adds Gaussian noise $\epsilon\sim N(0,1)$ to the original data that can be expressed as $q(x_{k}|x_{k-1})=N(\sqrt{1-\beta_{k}}x_{k-1},\beta_{k}\textbf{I})$, $\{\beta_{k}\in(0,1),k\in(1,K)\}$ is a set of scheduled noise weight. Based on the forward transition probability, the generated noised data in step k can be calculated by $x_{k}=\sqrt{\hat{\alpha}_{k}}x_{0}+(1-\hat{\alpha}_{k})\epsilon.$ The reversed chain utilizes a well-trained network $p_{\theta}$ to recurrently denoise $x_{K}$ for recovering original data $x_{0}$, which can be expressed as $p_{\theta}(x_{k-1}|x_{k})=N(\mu_{\theta}(x^{k},k),\sigma_{\theta}(x_{k},k)\textbf{I})$.
Training objective. It is essentially to maximize the log-likelihood function of the denoising network $p_{\theta}$ for the initial data $x_{0}$, i.e., $L(\theta)=\mathbb{E}_{x_{0}\sim q(x_{0})}\{-logp_{\theta}(x_{0})\}$. Subsequently, this function is optimized using the Variational Lower Bound (VLB), which can be expressed as:
Taking the expectation on both sides of the above equation and applying Fubini’s theorem (Yang et al., 2023b), we can derive:
(2) $\mathbb{E}_{q(x_{0})}\{-logp_{\theta}(x_{0})\}\leq\mathbb{E}_{q(x_{0})}\bigg{\{}\mathbb{E}_{q_{(}x_{1:T}|x_{0})}\{log\frac{q_{(}x_{1:T}|x_{0})}{p_{\theta}(x_{0:T}|x_{0}))}\}\bigg{\}}$
We can minimize the upper bound of $L(\theta)$ by minimizing $L_{vb}$, thereby maximizing the log-likelihood function of $p_{\theta}$. Ho et al. (Ho et al., 2020) proved that $L_{vb}(\theta)$ can be further parameterized by $\mu_{\theta}(x_{k},k)=\alpha_{k}^{-0.5}[x_{k}-\beta_{k}(1-\hat{\alpha}_{k})^{-0.5}\epsilon_{\theta}(x_{k},k)]$, and $\sigma_{\theta}$ can be parameterized as $\sigma_{\theta}(x_{k},k)=\sqrt{(1-\hat{\alpha}_{k-1})/(1-\hat{\alpha}_{k})\beta_{k}}$. The network $p_{\theta}$ can then be optimized by the following objective:
@@ 2.1. Diffusion model and training object
2.2. Problem formulation
Mobile Traffic refers to the volume of data transmitted over wireless channels between mobile devices and BSs within a period of time. In this paper, we primarily focus on the aggregated downlink mobile traffic generated by all mobile devices within the coverage area of BSs. We consider a discrete-time scenario $T=\{1,2,...t\}$ with equal time intervals. For a single BS, the traffic variation over time $T$ can be represented as $\{b_{t}\}_{t=1:T}$, where $b_{t}$ denotes the aggregated traffic within the coverage area of BS at time $t$. To characterize the mobile traffic features across an urban region $\mathcal{G}$, we define the geographical length and width of that region as $H$ and $V$, respectively. The mobile traffic of region $\mathcal{G}$ can then be defined as the sum of the aggregate traffic of all BSs located at $\mathcal{G}$: $\mathcal{S}_{t=1:T}=\sum_{b}\{b_{t}\}_{t=1:T}$. Based on the definition above, the mobile traffic forecasting problem can be viewed as predicting dynamics in future traffic $\mathcal{S}_{t=T:T_{0}}$, based on historical traffic information $\mathcal{S}_{t=1:T}$ and complex factors that influence network usage $C_{\mathcal{G}}$, such as user behavior and geographical contexts, which can be expressed as:
Probelm Definition: Given arbitrary urban region $\mathcal{G}$ with geographical length and width $H\times V$. The goal is to use a model $\mathcal{F}$ to predict the mobile traffic sequence $\mathcal{S}_{t=T:T_{0}}$ for the next $T_{0}$ time steps, conditioning on historical mobile traffic $\mathcal{S}_{t=1:T}$ and urban contextual factors $C$, i.e., $\mathcal{F}(\mathcal{S}_{t=T:T_{0}}|\mathcal{S}_{t=1:T},C_{\mathcal{G}})$.
There are two key requirements for achieving the above problem in large-scale urban environments. First, the model must be capable of processing massive traffic data across diverse urban spatiotemporal environments. Second, the model must flexibly handle various types of forecasting tasks in mobile networks. We identify 3 primary tasks:
$\bullet$ Short-term prediction task uses long historical data to predict mobile traffic dynamics over a short future period, $i.e.,$$\{\mathcal{S}_{t=1:T},C_{\mathcal{G}}\}\rightarrow\mathcal{S}_{t=T:T_{0}}$, where $T\geq(T_{0}-T)$. It mainly focuses on the fluctuations with less periodicity of traffic. Based on the forecasts, operators can understand upcoming network demands and make strategies for wireless resource and access control to improve user experience in real time.
$\bullet$ Long-term prediction task estimates future traffic patterns based on limited historical data, $i.e.,$$\{\mathcal{S}_{t=1:T},C_{\mathcal{G}}\}\rightarrow\mathcal{S}_{t=T:T_{0}}$, where $T\leq(T_{0}-T)$. It primarily explores inherent patterns and periodical features within the traffic. This forecasting enables operators to assess and analyze network performance from a global perspective, thereby facilitating the formulation of network optimization planning strategies, such as cell dormancy and network capacity expansion.
$\bullet$ Generation task focuses on identifying underlying network demand within a specific area without referring to historical data, $i.e.,$$\{C_{\mathcal{G}}\}\rightarrow\mathcal{S}_{t=1:T_{0}}$. Traffic generation helps operators assess potential communication demands in new regions lacking historical data, allowing them to develop planning strategies such as BS deployment, network segmentation, and capacity planning, etc.
However, building such a foundation model is not straightforward. Specifically, three key challenges arise: i). How can we generalize the large-scale traffic data across multiple cities and accurately model their rich spatiotemporal features? ii). What strategies can be employed for the training process, ensuring that the model is capable of handling the diverse forecasting tasks? iii). How to effectively integrate user dynamics and contextual features with mobile traffic?
3. Design of FoMo
3.1. Framework overview
Figure 1. The flowchart of the FoMo framework
To tackle the challenges, we propose the FoMo framework that incorporates three stages: data tokenization, model pre-training, and model fine-tuning, as illustrated in Figure 1.
i). Data tokenization. The stage reshapes mobile traffic data from various spatial ranges and time spans across multiple cities into a unified mobile token for model training and capturing the diverse features in mobile traffic.
ii). Masked diffusion-based pre-training. The stage enables the model to fully grasp the fundamental spatio-temporal features of mobile traffic across various forecasting tasks, where we design a masked diffusion model as the backbone, enabling the multitasking learning process.
iii). Urban context-aware fine-tuning. We design a contrastive learning algorithm that integrates external factors closely associated with mobile traffic, including network user dynamics and urban POI distributions.
3.2. Mobile traffic data tokenization
We draw inspiration from NLP tokenization (Dosovitskiy et al., 2021), where we decompose heterogeneous traffic data, with varying sampling intervals and diverse spatial ranges, into basic units. Embedding layers are then used to extract the intricate features of these tokens and learn their dependencies. Specifically, we define the minimum granularity of mobile tokens in space and time as $h_{0}$, $v_{0}$, and $t_{0}$, respectively. For traffic data $S$ of length $T$ within an urban region $H\times V$, the tokenization process breaks down $S$ into multiple small mobile tokens $X$, which can be expressed as
(4) $S\in\mathbb{R}^{H\times V\times T}\rightarrow X\in\mathbb{R}^{(H^{\prime}\times V^{\prime}\times T^{\prime})\times(h_{0}\times v_{0}\times t_{0})},$
where $H^{\prime}=H/h_{0}$, $V^{\prime}=V/v_{0}$, and $T^{\prime}=T/t_{0}$, the $(h_{0},t_{0},v_{0})$ of X represents the mobile token. Subsequently, we use an embedding layer $E_{x}(X)$ ($e.g.,$ pooling layer, convolutional layer, or fully connected layer) to map the mobile token with hidden features $C$, $i.e.,$ $E_{x}(X)\in\mathbb{R}^{(H^{\prime}\times V^{\prime}\times T^{\prime})\times C}$.
3.3. Masked diffusion-based pre-training
Pre-training primarily serves to enhance the foundation model’s understanding of various forecasting tasks and to capture the spatio-temporal correlations inherent in massive mobile data. We propose a masked diffusion model with self-supervised training (Zhao et al., 2024; He et al., 2022), where specific masks are tailored for the three forecasting tasks, as shown in Figure 2.
3.3.1. Multitasking mask
We develop 4 distinct masks $m\in\mathbb{R}^{H^{\prime}\times V^{\prime}\times T^{\prime}}$: short-term masking, long-term masking, generative masking, and random masking. The first three schemes align with specific forecasting tasks, while random masking aims to further explore spatio-temporal correlations within mobile traffic, and enhance its generalization capability.
$\bullet$ Short/Long-masking masking. The schemes mask the time dimension $T^{\prime}$ at a specific spatial location $(h,v)$, primarily to reconstruct the mobile traffic within the period $T^{\prime}-t_{0}$, where $t_{0}\in T^{\prime}$. Depending on the ratio of $t_{0}$ to $T^{\prime}$, the schemes correspond to short-term and long-term predictions, respectively, and can be expressed as:
@@ 3.3.1. Multitasking mask
$\bullet$ Generation masking. This type of mask fully obscures the temporal dimension of adjacent spatial regions $(h_{i\sim j},v_{i\sim j})$, enabling the model to generate complete temporal mobile traffic within the regions, where $i\sim j$ represents multiple consecutive adjacent mobile tokens. Unlike prediction masks that depend on prior historical temporal information, this mask requires capturing the spatio-temporal dependencies between the target spatial region and its surrounding regions to generate underlying distributions, which yields:
@@
$\bullet$ Random masking. The scheme masks mobile traffic across both spatial and temporal dimensions, which aims to capture diverse correlations of mobile tokens, aiding the model in understanding the complex features of mobile data. Denote $\mathcal{R}(H^{\prime},V^{\prime},T^{\prime})$ as randomly choosing items from $H^{\prime}$, $V^{\prime}$, and $T^{\prime}$, the masking scheme follows:

### v3

# UoMo: A Universal Model of Mobile Traffic Forecasting for Wireless Network Optimization

@@ Abstract.
Mobile traffic forecasting allows operators to anticipate network dynamics and performance in advance, offering substantial potential for enhancing service quality and improving user experience. It involves multiple tasks, including long-term prediction, short-term prediction, and generation tasks that do not rely on historical data. By leveraging the different types of mobile network data generated from these tasks, operators can perform a variety of network optimizations and planning activities, such as base station (BS) deployment, resource allocation, energy optimization, etc. However, existing models are often designed for specific tasks and trained with specialized data, and there is a lack of universal models for traffic forecasting across different urban environments. In this paper, we propose a Universal model for Mobile traffic forecasting (UoMo), aiming to handle diverse forecasting tasks of short/long-term predictions and distribution generation across multiple cities to support network planning and optimization. UoMo combines diffusion models and transformers, where various spatio-temporal masks are proposed to enable UoMo to learn intrinsic features of different tasks, and a contrastive learning strategy is developed to capture the correlations between mobile traffic and urban contexts, thereby improving its transfer learning capability. Extensive evaluations on 9 real-world datasets demonstrate that UoMo outperforms current models in various forecasting tasks and zero/few-shot learning. It shows an average accuracy improvement of 27.85%, 18.57%, and 15.6% in long-term prediction, short-term prediction, and generation tasks, respectively, showcasing its strong forecasting capability. We deploy UoMo on China Mobile’s JiuTian platform, leveraging the predicted mobile data to optimize live networks. This optimization includes BS deployment, resulting in a 25.3% increase in served users, and BS sleep control, which reduces equipment depreciation by 40.7%. The source code is available online: https://github.com/tsinghua-fib-lab/UoMo.
Mobile traffic forecasting; foundation model; diffusion model
journalyear: 2025conference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2; August 3–7, 2025; Toronto, ON, Canadabooktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadadoi: 10.1145/3711896.3737272isbn: 979-8-4007-1454-2/2025/08ccs: Computing methodologies Knowledge representation and reasoningccs: Networks Network simulationsccs: Information systems Spatial-temporal systems
@@ 1. Introduction
In recent years, foundation models (brown2020language, 6, 48, 42) have made substantial strides in natural language processing and computer vision. These models are reshaping the AI ecosystem by harnessing their powerful data processing, generalization, and zero/few-shot learning capabilities. An increasing number of specialized domains have developed foundational models tailored to their specific data and contextual demands, including healthcare, medicine, urban navigation, and beyond (10.1145/3636534.3649361, 68, 65, 24, 4). Mobile networks encompass massive amounts of mobile traffic, user, and geographical data, providing inherent data support for building universal models. However, such dedicated models for mobile network domains have yet to be established. We hence aim to construct a universal model for mobile traffic forecasting, which can handle rich features of large-scale mobile data while retaining the generalization needed across multiple network applications (10.1109/TNET.2021.3136707, 54, 8, 67).
@@
Constructing such a universal model of mobile traffic forecasting is vital for mobile networks (10.1145/3637528.3671730, 14, 53, 76). On the one hand, mobile traffic forecasting offers great potential for network planning and optimization. It enables operators to anticipate traffic dynamics, facilitating proactive perceptions of resource utilization and service quality, and allowing for the preemptive development of optimization strategies. On the other hand, mobile networks encompass a variety of optimization scenarios, including radio resource scheduling (6168145, 36, 32), Base Station (BS) deployment (9204664, 12, 13), and antenna configuration (10086045, 34, 29), etc. These scenarios involve diverse objectives like throughput, coverage, and energy efficiency, which impose distinct tasks on traffic forecasting. For example, radio resource scheduling requires performing short-term traffic forecasting tasks, prioritizing traffic dynamics to improve user experience (8896929, 9, 16), whereas BS deployment involves long-term traffic forecasting tasks, focusing on long-term traffic patterns within a region to align with network demands (9586045, 75, 52). For the planning and deployment of live wireless networks, it is essential to leverage the powerful data mining and robust generalization capabilities of universal models to simultaneously address a variety of optimization tasks.
@@
Although numerous notable works have emerged in the area of mobile traffic forecasting (10.1145/3589132.3625569, 17, 25, 39, 26, 58, 71, 45), current methods typically employ one-to-one approaches: designing customized models by leveraging task-specific data (8641378, 63, 20, 76, 28). The complex, customized models deployed in live networks often require manual orchestration and scheduling, which will lead to the waste of computational and storage resources, increasing the overhead of model deployment. In addition, mobile traffic is inherently heterogeneous of various collection granularity and scope. For example, Measure Report (MR) data primarily collects millisecond-level user traffic, while Performance Management (PM) data gathers cell-level traffic statistics over 15-minute intervals (7299258, 37, 51), leading to the absence of a unified representation akin to that found in natural language. Consequently, it is challenging to directly apply pre-trained models from the natural language/visual domains to mobile traffic data. Although some efforts have been made to reprogram mobile traffic data into a natural language format (NEURIPS2023_3eb7ca52, 18, 23), this approach heavily relies on the quality of manually crafted prompts, making it difficult to capture a universal representation of mobile traffic. Specifically, current mobile traffic forecasting models face two key limitations:
i) Limited generalization. Mobile traffic data is inherently shaped by the spatio-temporal dynamics of population distribution and communication demands. Due to variations in geographic environments, lifestyle habits, and urban layouts across different cities, mobile traffic can differ significantly (8984314, 66, 57). With relatively small parameters, current models struggle to capture the diverse spatio-temporal patterns inherent in large-scale data across multiple cities. Additionally, it is challenging to encapsulate the complex correlations between contextual factors and mobile traffic, resulting in poor transferability in multi-city scenarios.
ii) Constrained task adaptability. Mobile traffic forecasting is extensively applied across varying optimization scenarios. However, current models are often designed with specialized modules tailored to specific tasks. For instance, in short-term forecasting, models usually focus on capturing traffic fluctuations that employ autoregressive or event-driven methods. In contrast, long-term predictions emphasize the regular patterns of traffic and typically utilize time series decomposition techniques. These dedicated models increase design complexity and raise deployment costs when applied to diverse scenarios.
@@
To tackle the limitations, we propose a Universal model for Mobile traffic forecasting (UoMo), which aims to learn universal features of mobile traffic data and to handle multiple tasks in mobile networks, thereby establishing a one-for-all forecasting model. First, inspired by Sora (videoworldsimulators2024, 5), UoMo adopts the transformer-based diffusion model as the backbone instead of the U-Net structure, to help the model understand the diverse features of massive mobile data. We propose a contrastive diffusion algorithm and adjust the variational lower bound by analyzing the cross-entropy between mobile traffic and contextual features. This helps the model better integrate environmental information, improving generalization and addressing the first limitation. Second, we adopt a task-oriented masking and self-supervised training paradigm, where we categorize traffic forecasting in mobile networks into three tasks: short-term prediction, long-term prediction, and generation. We design the corresponding masking strategies to enable the model to learn data features for various tasks and adapt to multiple tasks, thus addressing the second research challenge.
@@
$\bullet$ To the best of our knowledge, it is the first universal model designed for mobile traffic forecasting. The proposed model enables various forecasting tasks in mobile networks across different urban environments via a unified framework, assisting network operators in achieving highly efficient network planning and optimization.
@@
$\bullet$ We develop our universal model using a masked diffusion approach with spatio-temporal masking strategies tailored for diverse forecasting tasks, including short/long-term predictions and distribution generation. To strengthen the correlation between contextual features and mobile traffic, we further propose a context-aware contrastive learning fine-tuning strategy, which can enhance forecasting and transfer learning capabilities.
$\bullet$ We conduct extensive evaluations with 9 real-world mobile traffic datasets. The results validate UoMo’s superior generalization, multi-task capabilities, and robust few/zero-shot performance in unseen scenarios. We also deploy the UoMo model on China Mobile’s Jiutian platform. Experiments on the live system data prove our model empowers multiple network optimization scenarios, including a 25.3% increase in served users for BS deployment and a 40.7% reduction in equipment depreciation for BS sleep control.
2. Problem Formulation
Mobile Traffic refers to the volume of data transmitted over wireless channels between mobile devices and BSs over a period of time. We consider a discrete-time scenario $T=\{0,...T\}$ with equal time intervals. For a single BS, the traffic variation over time $T$ can be represented as $\{b_{t}\}_{t=0:T}$, where $b_{t}$ denotes the aggregated traffic within the coverage area of BS at time $t$. To characterize the mobile traffic features across an urban region $\mathcal{G}$ with contextual features $C_{\mathcal{G}}$, we define the geographical length and width of that region as $H$ and $V$, respectively. The mobile traffic of region $\mathcal{G}$ can then be defined as the sum of the aggregate traffic of all BSs located at $\mathcal{G}$: $\mathcal{S}_{0:T}=\sum_{b}\{b_{t}\}_{t=0:T}$. Regarding diverse tasks in communication networks, such as base station deployment, user access control, and wireless resource allocation, different traffic forecasting tasks are often required. We categorize forecasting into 3 typical tasks:
$\bullet$ Short-term prediction task uses long historical data $0\sim t$ to predict mobile traffic dynamics over a short future period, $i.e.,$$\{\mathcal{S}_{0:t},C_{\mathcal{G}}\}\rightarrow\mathcal{S}_{t:T}$, where $t\gg(T-t)$. Based on the forecasts with fluctuations, operators can understand upcoming network demands and make optimizations such as resource allocation (9923420, 35) and access control (10255286, 31) to improve user experience in live networks.
$\bullet$ Long-term prediction task estimates future traffic patterns based on limited historical data, $i.e.,$$\{\mathcal{S}_{0:t},C_{\mathcal{G}}\}\rightarrow\mathcal{S}_{t:T}$, where $t\ll T$. It explores inherent periodical patterns within the traffic. This forecasting enables operators to estimate and analyze network performance from a global perspective, thereby facilitating the formulation of network optimization planning strategies, such as cell dormancy (10390348, 56) and network capacity expansion (10063388, 78).
$\bullet$ Generation task focuses on identifying underlying network demand within a specific area without referring to historical data, $i.e.,$$\{C_{\mathcal{G}}\}\rightarrow\mathcal{S}_{t=0:T}$. It helps operators assess potential communication demands in new regions lacking historical data, allowing them to develop planning strategies such as BS deployment (10086045, 34), network segmentation (10634862, 60), and capacity planning (10164177, 33), etc.
We aim to build a domain-universal model capable of achieving the above 3 forecast tasks. The problem can be defined as follows.
Problem Definition. Given an arbitrary urban region $\mathcal{G}$, the goal is to use a model $\mathcal{F}$ to forecast diverse mobile traffic sequence $\mathcal{S}$ with short/long/generation tasks, conditioning on urban contextual factors $C_{\mathcal{G}}$, i.e., $\mathcal{F}(\mathcal{S}_{t=T_{0}:T_{1}}/\mathcal{S}_{t=0:T_{1}},C_{\mathcal{G}})$.
However, building such a universal model is not straightforward. Specifically, two key challenges arise: i). What strategies can be employed for the training process, ensuring that the model can handle the diverse forecasting tasks? ii). How to effectively integrate user dynamics and contextual features with mobile traffic?
3. Design of UoMo
3.1. Framework overview
Figure 1. The flowchart of the UoMo framework
To tackle the challenges, we propose the UoMo framework incorporating three stages as illustrated in Figure 1.
i). Data tokenization reshapes mobile traffic data from various spatial-temporal spans across multiple cities into a unified mobile token for model training and capturing their diverse features.
ii). Masked diffusion-based pre-training tends to fully grasp the fundamental spatio-temporal features of mobile traffic across various forecasting tasks, where we design a diffusion-based backbone and task-oriented masks.
iii). Urban context-aware fine-tuning introduces a contrastive learning algorithm that integrates external factors closely associated with mobile traffic, including network user dynamics and urban POI distributions.
3.2. Masked diffusion-based pre-training
We propose a masked diffusion model with self-supervised training, where specific masks are tailored for the three forecasting tasks to enhance the model’s understanding of various forecasting tasks and to capture the spatio-temporal correlations inherent in massive mobile data, as shown in Figure 2.
3.2.1. Mobile traffic data tokenization
We draw inspiration from NLP tokenization, where we decompose traffic data with varying sampling intervals and diverse spatial ranges into basic unit $h_{0}\times v_{0}\times t_{0}$. For traffic data $S$ of length $T$ within an urban region $H\times V$, the tokenization process breaks down $S$ into multiple small mobile tokens $X$, which can be expressed as $S\in\mathbb{R}^{H\times V\times T}\rightarrow X\in\mathbb{R}^{(H^{\prime}\times V^{\prime}\times T^{\prime})\times(h_{0}\times v_{0}\times t_{0})},$ where $H^{\prime}=H/h_{0}$, $V^{\prime}=V/v_{0}$, and $T^{\prime}=T/t_{0}$, the $(h_{0},t_{0},v_{0})$ of X represents the mobile token. Subsequently, we use an embedding layer $E_{x}(X)$ ($e.g.,$ pooling layer, convolutional layer, or fully connected layer) to map the mobile token with hidden features $C$, $i.e.,$ $E_{x}(X)\in\mathbb{R}^{(H^{\prime}\times V^{\prime}\times T^{\prime})\times C}$.
3.2.2. Task-oriented mask
After the mobile traffic data tokenization, the original $H\times V$ region is divided into multiple $h\texttimes v$ areas. The masking strategy performs self-supervised training by masking and reconstructing partial areas. We define the masked parts as the target areas, and the unmasked parts around the target areas as the surrounding areas. We develop 4 distinct masks: short-term, long-term, generation, and random masks, $m\in\mathbb{R}^{H^{\prime}\times V^{\prime}\times T^{\prime}}$. The first three focus on specific forecasting tasks, while random masking explores spatio-temporal correlations to enhance generalization.
@@ 3.2.2. Task-oriented mask
$\bullet$ Generation mask. The generation mask completely obscures the temporal dimension at a specific spatial location $(h,v)$, enabling the model to generate mobile traffic sequence within the target area. Unlike prediction masks that rely on historical data, it captures spatio-temporal dependencies between the target area and its surrounding areas to generate underlying distributions:
@@
$\bullet$ Random mask. It masks mobile traffic across both spatial and temporal dimensions, which aims to capture diverse correlations of mobile tokens, aiding the model in understanding the complex features of mobile data. Denote $\mathcal{R}(H^{\prime},V^{\prime},T^{\prime})$ as randomly choosing items from $H^{\prime}$, $V^{\prime}$, and $T^{\prime}$: